OpenStudy (anonymous):

Help me please :D Design at least three different equivalent forms of x^2 Show how each one can be simplified back to x^2 in two or more steps HELP Please!! :)

OpenStudy (anonymous):

x*x=x^2

OpenStudy (anonymous):

like thAT

ganeshie8 (ganeshie8):

one form that comes to my mind is \(\large \sqrt{x^4}\)

ganeshie8 (ganeshie8):

second form \(\large \frac{x^5}{x^3} \) here \(x \ne 0\)

OpenStudy (anonymous):

can't use x*x @AllTheLonelyKillers because that is only one step.. :)

ganeshie8 (ganeshie8):

can u think of a third form ? maybe try putting x^2 inside log or somthing..

OpenStudy (anonymous):

what is log? @ganeshie8

OpenStudy (anonymous):

oh i was just giving a guess i am kinda clueless here

ganeshie8 (ganeshie8):

forget it, think of someother rational funciton like secondone may be :)

OpenStudy (anonymous):

Well thank you :) @AllTheLonelyKillers

OpenStudy (anonymous):

so I just put any number with the x & solve it? @ganeshie8

ganeshie8 (ganeshie8):

third form :- \(\large \frac{x^3(x-1)}{x^2-x}\)

ganeshie8 (ganeshie8):

if you simplify above 3 forms, you wud end up wid \(x^2\)

ganeshie8 (ganeshie8):

but, each has its own domain restrictions. (ur teacher wants u to think about domain restrictions)

OpenStudy (anonymous):

what kind of domain restrictions?

ganeshie8 (ganeshie8):

before we get that that, are u convinced, all above forms simplify to \(x^2\) ? :)

OpenStudy (anonymous):

I believe you lol b/c I have no clue how to do any of that

ganeshie8 (ganeshie8):

it is good if u r able to simplify them and see that they indeed give u x^2, let me show u how to simplify each of them quick

ganeshie8 (ganeshie8):

first form :- \(\large \sqrt{x^4} = \sqrt{(x^2)^2} = \color{red}{x^2}\) second form :- \(\large \frac{x^5}{x^3} = \frac{x^2x^3}{x^3} = \color{red}{x^2} \) third form :- \(\large \frac{x^3(x-1)}{x^2-x} = \frac{x^2x(x-1)}{x(x-1)} = \color{red}{x^2}\)

ganeshie8 (ganeshie8):

see if they make sense domain restrictions, id let ur teacher explain u... it will be next topic maybe :) good luck !

OpenStudy (anonymous):

you're amazing! thank you :)

ganeshie8 (ganeshie8):

aww you're sweet, yu wlcme :D